OpenStudy (anonymous):

Write an exponential function y=ab^x for a graph that will include that will include (0,4) and (2, 49).

OpenStudy (anonymous):

I am not sure how to write equations, and I could really use the help on this.

OpenStudy (reemii):

\(ab^0 = ?\)

OpenStudy (reemii):

\[ \begin{cases} ab^0 = a & \\ ab^0 = 4 &\text{ as asked} \end{cases} \] therefore \(a=4\). Now you can find \(b\).

OpenStudy (kenljw):

4=a b^0 b^0 =1 a=4 49 = 4 b^2 b = (1/2)log(49/4)

OpenStudy (kenljw):

Principle log(x^n) = n log(x)